# RDA-PSO: A computational method to quantify the diffusive dispersal of insects

### v1

@@ Abstract
This article introduces a computational method, called Recapture of Diffusive Agents & Particle Swarm Optimization (RDA-PSO), designed to estimate the dispersal parameter of diffusive insects in mark-release-recapture (MRR) field experiments. In addition to describing the method, its properties are discussed, with particular focus on robustness in estimating the observed diffusion coefficient. It is shown that RDA-PSO provides a simple and reliable approach to quantify mosquito dispersal that outperforms other techniques based on the solution of the diffusion equation, three of which are also introduced in this work. Examples of application to both synthetic and real field data for the yellow fever mosquito are provided.
@@
Diffusion, Estimation of dispersal properties, Particle Swarm Optimization, Mark-release-recapture experiments, Mosquitoes, Parameter inference
@@ 1 Introduction
This article focuses on diffusive models of Ae. aegypti dispersal and presents methods to measure the associated diffusion coefficient. In particular, we introduce a computational model that accurately estimates diffusive dispersal parameters associated with mark-release-recapture (MRR) data, using particle swarm optimization. In MRR experiments, a number of marked insects are released in the environment at the same time from a single location. In the following days and weeks, researchers systematically search for the released insects at regular times and record when and where they were recaptured. Typically, only a few percent of the released insects are found. We analyze the robustness of the computational approach presented here with respect to changes in experimental parameters (e.g. size and attractiveness of insect collection sites) and quantify the uncertainty associated with both experimental conditions and computational limitations. Section 2 reviews the literature on mosquito dispersal, focusing on why a diffusive process (possibly with advection in the presence of wind or bias towards attracting sites) is a reasonable approximation. Section 3 introduces a simple calculation based on the mean distance traveled (MDT) by mosquitoes recaptured in MRR experiments, as well as three optimization-based approaches to estimating the diffusion coefficient from recapture data. The time-corrected and area-and-time-corrected models use theoretical expressions derived from the diffusion equation and are based on the assumption that recapturing released mosquitoes (often done by aspirating insects found in a specific location at a given time or by using traps) amounts to sampling their overall distribution. The Recapture of Diffusive Agents & Particle Swarm Optimization (RDA-PSO) model, which is the main focus of this article, simulates insect diffusion and recapture at the microscopic level. Section 4 discusses the performance of the three optimization-based models on synthetic and real-world recapture data and numerically quantifies uncertainty associated with RDA-PSO estimates. Section 5 summarizes the results and highlights the superiority of the proposed computational approach in estimating dispersal parameters from noisy data. In particular, the RDA-PSO method introduced here is robust to lack of knowledge on the precise location of collection sites, its input data do not have to be corrected to account for unevenly distributed recapture locations, and it works well even when the number of recaptured mosquitoes is small, which is typical of mark-release-recapture experiments.
@@ 2 Modeling of mosquito dispersal
Mosquitoes typically disperse to rest, mate, feed (on nectar or on blood), and to lay eggs [62]. Early quantitative studies of their dispersal behavior go back more than 100 years (Ross’s work [57, 58] cited by Service [62]). Mark-release-recapture (MRR) experiments, in which tagged insects are released in the environment and recaptured at later times, have been used to measure native population abundance or compare survival rates of different morphs in a polymorphic population (see e.g. [45]). They also provide a way to quantify arthropod dispersal. Typically, concentric annular zones are defined about the release point and numbers of marked insects captured in each zone are reported during a period of a few days, weeks or even months after the initial release. Recapture may be done by traps or by systematic aspiration in pre-specified areas (e.g. houses). For more details, the reader is referred to [65]. Two concepts are central to the analysis of MRR data. First, one needs to model how mosquitoes move; we will see that there is plenty of evidence to describe such dynamics in terms of regular diffusion. Second, one needs to assess how to relate recapture data to mosquito density.
@@ 2.1 Dispersal at the microscopic level
It is natural to describe dispersal at the level of individual insects in terms of a random walk [49, 8], consisting of successive displacements along straight lines alternating with changes in the direction (or angle with respect to an axis of reference) in which each straight forward motion occurs. The use of random walks in biology is of course very common [4, 8], from modeling how bacteria swim by alternating runs and tumbles [4] to, at a much bigger scale, how animals look for food. In many situations, these walks are expected to be correlated [31, 1], due to a general tendency of moving forward [8]: each new direction of motion is close to the preceding one, thereby limiting the likelihood of large turning angles. However, after many steps, such a random process is equivalent to an isotropic random walk, in the sense that the mean square displacement scales linearly with the number of steps [31, 53, 1]. In other words, the memory of short-term correlations is eventually lost [8] and, in the absence of overall bias, insect dispersal is well described by isotropic diffusion at the macroscopic level [8, 53, 1].
@@ 2.2 Dispersal at the macroscopic level
2.3 Mean distance traveled
@@ 2.5 Computational models of mosquito dispersion
Diffusion-based computational models of mosquito dispersal are, broadly speaking, agent-based, PDE-based, or discrete meta-population models. Agent-based models [54, 84, 26, 67, 81, 44, 13] describe dynamics at the microscopic level, by following individual insects that perform isotropic or correlated random walks. Some meta-population models such as [51, 40, 35, 50, 23, 11, 10] describe movement in a heterogeneous landscape in terms of flux or transfer rate between different patches, including regular diffusion within a patch [40, 35] or long-range dispersal due to transportation networks [11, 10]. Other applications of diffusive behavior include the use of time-dependent diffusion kernels to assess the likelihood of mosquitoes moving between unsealed water tanks in an urban environment [74].
In the following, we introduce theoretical and computational methods to estimate the diffusive properties of dispersing insects from MRR experiments. By analyzing the performance of these methods against synthetic data, for which the actual diffusion coefficient is known, we conclude that approaches based on the solution of the diffusion equation, although theoretically correct, are in practice inaccurate, whereas the RDA-PSO method produces robust results that do not require additional corrections. Consequently, many quantitative estimates related to mosquito dispersal found in the MRR literature may need to be revisited. Although diffusive processes are well understood, the novelty of our contribution resides in its ability to address small recapture rates, the non-uniform placement of recapture locations, and the uncertainty related to how mosquitoes are attracted to capture locations.
@@ 3 Methodology
This section introduces four methodological approaches designed to estimate the diffusion coefficient associated with MRR data. Theoretical considerations are based on the solution of the diffusion equation and serve as baselines against which the computational model is compared. The latter model, which we argue is better fitted to analyze MRR experiments, relies on a description of insect dynamics at the microscopic level. All methods assume that mosquitoes simultaneously and independently diffuse from a central release point before being recaptured by collection traps located in concentric zones of width $r_{zone}$ centered on that point. More precisely, each zone is an annulus of inner radius $r_{j-1}$ and outer radius $r_{j}$, where $r_{j}=j\,r_{zone}$, $j=1,2,\cdots,n_{z}$, and $n_{z}$ is the total number of zones (see for instance Figure 4 of Appendix D for an illustration). The surface area of each zone increases with its index $j$. Empirical recapture ratios, defined in Section 3.2, are used as input data.
3.1 Recapture probabilities associated with diffusive spread
@@
3.2 Empirical recapture ratios
$\tau_{i}=\frac{\text{number of mosquitoes recaptured on day $i$}}{\text{total number of mosquitoes recaptured}}.$ (4)
and spatial ratios are defined by
$\sigma_{j}=\frac{\text{number of mosquitoes recaptured in zone $j$}}{\text{total number of mosquitoes recaptured}},$ (5)
@@ 3.3 Time-corrected model
One of the main difficulties of translating MRR data into estimates of dispersal parameters is that only a very small percentage of released mosquitoes is eventually recaptured. To improve empirical estimates of the quantities defined in Section 3.1, it is therefore useful to combine information associated with different collection dates. To this end, we estimate the probability density function for a mosquito being present at a distance $r$ from the point of release over the entire recapture experiment,
@@
as a weighted average of the probabilities $P_{t}(r)$ defined in Equation (1), where the weights reflect the fraction of mosquitoes recaptured each day $i$. Thus, we set
[…]
where $N$ is the number of mosquitoes initially released and $K$ measures the total recapture ratio. However, in the presence of depletion, due to mosquitoes being caught or dying, the number captured on day $i$ is $KN_{i}$, where $N_{i}$ is close to but slightly less than $N$, which leads to $\tau_{i}=N_{i}/\left(\sum_{i=1}^{n}N_{i}\right)$. The definition of $P_{tot}$ is such that
[…]
@@
Estimates of $Q_{tot}$ (and $P_{tot}$) may be compared directly to data observed in MRR experiments, and the fit between model and experiment may be optimized by selecting an appropriate value of the diffusion coefficient $D$. We therefore define the time-corrected model,
@@ 3.4 Area-and-time-corrected model
where $N_{j}$ is the number of mosquitoes actually captured in zone $j$, and $nT_{j}\neq 0$ is the number of traps placed in zone $j$. The above equation can be recast as $N_{j}^{c}=N_{j}\cdot\rho_{tot}/\rho_{j}$, with $\rho$ denoting the density of traps per region, which shows how the correction compensates for uneven trap density. Such a correction changes the number of mosquitoes recaptured in each zone every day of the experiment and thus affects the spatial and temporal ratios defined in Equations (5) and (4). Fitting these corrected ratios into (9) leads to a corrected estimate of the diffusion coefficient, which we refer to as $D_{ATC}$.
3.5 Estimating the diffusion coefficient from the empirical MDT
It may be tempting to estimate the diffusion coefficient $D$ from the mean distance traveled (MDT) by recaptured mosquitoes, whose daily values are sometimes, but not always, reported in the literature. As already mentioned, estimating $D$ from such information is not easily done because of the low recapture rate. First, we note that a simple random walk simulation shows that estimating the diffusion coefficient from $\langle r^{2}\rangle_{t}$ is more reliable than using the MDT. Second, and more importantly, computer experiments require at least 1000 random walkers to produce a reasonable estimate of the diffusion coefficient from the MDT. Thus, results based on a few dozens of mosquitoes are unlikely to give reliable or consistent estimates from day to day. It is therefore essential to aggregate the results obtained over different days in order to improve the estimate, since data from a larger number of walkers are then used to arrive at a single number. To this end, we define the MDT of an entire experiment as a weighted average of the daily MDT’s, consistent with the first equation of (3) and in a way similar to Equations (6) and (7) above:
$MDT_{tot}$ $=\sum_{i=1}^{n}\tau_{i}\,\sqrt{\pi Dt_{i}}$
[…]
We emphasize that in Equation (11), the quantity $MDT_{tot}$ defines an overall MDT, calculated for all recaptured mosquitoes, regardless of the day of recapture. This quantity is expected to be close to the empirical MDT, $MDT_{emp}$, reported in the literature (e.g. as defined in [38]). Indeed,
$\simeq\sum_{i=1}^{n}\tau_{i}\,MDT_{i}=\sum_{i=1}^{n}\tau_{i}\sqrt{\pi Dt_{i}}=\sqrt{\pi D}\sum_{i=1}^{n}\tau_{i}\sqrt{t_{i}}=MDT_{tot}.$
In the above, $N_{tot}=\sum_{i=1}^{n}N_{i}$ is the total number of recaptured mosquitoes, $r_{k}$ is the distance from the release point where the $k$-th mosquito recaptured on day $i$ was caught, and $MDT_{i}$ is the MDT for day $i$. The difference between $MDT_{emp}$ and $MDT_{tot}$ is that the exact expression of $MDT_{i}$ is used in the latter, instead of its empirical approximation $\sum_{k=1}^{N_{i}}\frac{r_{k}}{N_{i}}$ used in the former. Knowledge of the empirical MDT leads to the following estimate of the diffusion coefficient,
$D_{MDT}=\frac{1}{\pi}\left(\frac{MDT_{emp}}{\sum_{i=1}^{n}\tau_{i}\sqrt{t}_{i}}\right)^{2}.$ (12)

### v2

@@ Abstract
This article introduces a computational method, called Recapture of Diffusive Agents & Particle Swarm Optimization (RDA-PSO), designed to estimate the dispersal parameter of diffusive insects in mark-release-recapture (MRR) field experiments. In addition to describing the method, its properties are discussed, with particular focus on robustness in estimating the observed diffusion coefficient in the presence of uncertainty. It is shown that RDA-PSO provides a simple and reliable approach to quantify insect dispersal that can handle low recapture rates and uneven capture site distributions without the need for area corrections. Tests on synthetic data, for which the actual diffusion coefficient is known, show the method outperforms three techniques based on the solution of the diffusion equation, which are also introduced in this work. Examples of application to real field data for the yellow fever mosquito are provided.
@@
Diffusion, Estimation of dispersal properties, Particle Swarm Optimization, Mark-release-recapture experiments, Mosquitoes, Parameter inferencze
@@ 1 Introduction
This article focuses on diffusive models of Ae. aegypti dispersal and presents methods to measure the associated diffusion coefficient. In particular, we introduce a computational model that accurately estimates diffusive dispersal parameters associated with mark-release-recapture (MRR) data, using particle swarm optimization. In MRR experiments, a number of marked insects are released in the environment at the same time from a single location. In the following days and weeks, researchers systematically search for the released insects at regular times and record when and where they were recaptured. Typically, only a few percent of the released insects are found. We analyze the robustness of the computational approach presented here with respect to changes in experimental parameters (e.g. size and attractiveness of insect capture sites) and quantify the uncertainty associated with both experimental conditions and computational limitations. Section 2 reviews the literature on mosquito dispersal, focusing on why a diffusive process (possibly with advection in the presence of wind or bias towards attracting sites) is a reasonable approximation. Section 3 introduces a simple calculation based on the mean distance traveled (MDT) by mosquitoes recaptured in MRR experiments, as well as three optimization-based approaches to estimating the diffusion coefficient from recapture data. The time-corrected and area-and-time-corrected models use theoretical expressions derived from the diffusion equation and are based on the assumption that recapturing released mosquitoes (often done by aspirating insects found in a specific location at a given time or by using traps) amounts to sampling their overall distribution. The Recapture of Diffusive Agents & Particle Swarm Optimization (RDA-PSO) model, which is the main focus of this article, simulates insect diffusion and recapture at the microscopic level. Section 4 discusses the performance of the three optimization-based models on synthetic and real-world recapture data and numerically quantifies uncertainty associated with RDA-PSO estimates. Section 5 summarizes the results and highlights the strengths of the proposed computational approach in estimating dispersal parameters from noisy data. In particular, the RDA-PSO method introduced here is robust to lack of knowledge on the precise location of capture sites, its input data do not have to be corrected to account for unevenly distributed recapture locations, and it works well even when the number of recaptured mosquitoes is small, which is typical of mark-release-recapture experiments.
@@ 2 Modeling of mosquito dispersal
Mosquitoes typically disperse to rest, mate, feed (on nectar or on blood), and to lay eggs [66]. Early quantitative studies of their dispersal behavior go back more than 100 years (Ross’s work [61, 62] cited by Service [66]). Mark-release-recapture (MRR) experiments, in which tagged insects are released in the environment and recaptured at later times, have been used to measure native population abundance or compare survival rates of different morphs in a polymorphic population (see e.g. [48]). They also provide a way to quantify arthropod dispersal. Typically, concentric annular zones are defined about the release point and numbers of marked insects captured in each zone are reported during a period of a few days, weeks or even months after the initial release. Recapture may be done by traps or by systematic aspiration in pre-specified areas (e.g. houses). In this article, we use the words trap or capture site interchangeably. For more details, the reader is referred to [69]. Two concepts are central to the analysis of MRR data. First, one needs to model how mosquitoes move; we will see that there is plenty of evidence to describe such dynamics in terms of regular diffusion. Second, one needs to assess how to relate recapture data to mosquito density.
@@ 2.1 Dispersal at the microscopic level
It is natural to describe dispersal at the level of individual insects in terms of a random walk [53, 10], consisting of successive displacements along straight lines alternating with changes in the direction (or angle with respect to an axis of reference) in which each straight forward motion occurs. The use of random walks in biology is of course very common [5, 10], from modeling how bacteria swim by alternating runs and tumbles [5] to, at a much bigger scale, how animals look for food. In many situations, these walks are expected to be correlated [33, 2], due to a general tendency of moving forward [10]: each new direction of motion is close to the preceding one, thereby limiting the likelihood of large turning angles. However, after many steps, such a random process is equivalent to an isotropic random walk, in the sense that the mean square displacement scales linearly with the number of steps [33, 57, 2]. Thus any local directional bias arising from correlated consecutive steps diminishes over time, and the resulting long-term dynamics may be replaced by a regular random walk, with uniformly distributed step orientations. In other words, the memory of short-term correlations is eventually lost [10] and, in the absence of overall bias, insect dispersal is well described by isotropic diffusion at the macroscopic level [10, 57, 2].
@@ 2.2 Dispersal at the macroscopic level
2.3 Estimation of dispersal-related bionomic parameters
@@ 2.4 Computational models of mosquito dispersion
Diffusion-based computational models of mosquito dispersal are, broadly speaking, agent-based, PDE-based, or discrete meta-population models. Agent-based models [58, 88, 28, 71, 85, 47, 15] describe dynamics at the microscopic level, by following individual insects that perform isotropic or correlated random walks. Some meta-population models [55, 43, 37, 54, 25, 13, 12] describe movement in a heterogeneous landscape in terms of flux or transfer rate between different patches, including regular diffusion within a patch [43, 37] or long-range dispersal due to transportation networks [13, 12]. Other applications of diffusive behavior include the use of time-dependent diffusion kernels to assess the likelihood of mosquitoes moving between unsealed water tanks in an urban environment [78].
In the following, we introduce theoretical and computational methods to estimate the diffusive properties of dispersing insects from MRR experiments. They all use the same data as input, specifically the temporal and spatial ratios defined in Equations (4) and (5) below. By analyzing the performance of these specific methods against synthetic data, for which the actual diffusion coefficient is known, we conclude that the three approaches based on the solution of the diffusion equation, although theoretically correct, are in practice inaccurate, whereas the RDA-PSO method produces robust results that do not require additional corrections. Although diffusive processes are well understood, the novelty of our contribution resides in its ability to address small recapture rates, the non-uniform placement of recapture locations, and the uncertainty related to how mosquitoes are attracted to capture locations.
@@ 3 Methodology
Our goal is to estimate the diffusion coefficient associated with MRR data. As described in Section 2, it is reasonable to assume the motion of insects in MRR experiments is diffusive in the region of study, away from attracting zones associated with capture sites (houses or traps). We also assume that the experiments are conducted in the absence of wind strong enough to carry some of the insects away (in particular, Aedes aegypti does not typically fly at high altitudes, although it has been observed in upper floors of high-rise buildings [1]). Heterogeneity due to landscape features is restricted to those representing capture sites. We do not factor in explicit mortality rates into our assumptions, although the daily temporal ratios defined in Equation (4) allow for depletion of the number of tagged mosquitoes in the environment, for instance as a consequence of mortality or because they left the area of study.
In this section, we introduce four methodological approaches to calculate the diffusion coefficient: the empirical MDT-based estimate (section 3.2), the time-corrected model (section 3.3), the area-and-time-corrected model (section 3.4), and the RDA-PSO computational model (section 3.5). The first three stem from theoretical considerations that are based on the solution of the diffusion equation and serve as baselines against which the computational model is compared. The latter model, which we argue is better fitted to analyze MRR experiments, relies on a description of insect dynamics at the microscopic level. All methods require the same set of input data, which we present next.
3.1 Preliminaries and input data
We refer to the region of study as $\mathcal{R}$, and assume that mosquitoes simultaneously and independently diffuse from a central release point before being recaptured by traps located in concentric zones of width $r_{zone}$ centered on that point. More precisely, each zone in $\mathcal{R}$ is an annulus of inner radius $r_{j-1}$ and outer radius $r_{j}$, where $r_{j}=j\,r_{zone}$, $j=1,2,\cdots,n_{z}$, and $n_{z}$ is the total number of zones (see for instance Figure 4 of Appendix D for an illustration). The surface area of each zone increases with its index $j$. Empirical recapture ratios, defined in Section 3.1.2, are used as input data.
3.1.1 Recapture probabilities associated with diffusive spread
@@
$Q_{t}(r_{a},r_{b})=\int_{r_{a}}^{r_{b}}P_{t}(r)\,dr=\exp\left(-\frac{r_{a}^{2}}{4Dt}\right)-\exp\left(-\frac{r_{b}^{2}}{4Dt}\right).$ (2)
$\langle r\rangle_{t}=\int_{0}^{\infty}rP_{t}(r)\,dr=\sqrt{\pi Dt},\qquad\langle r^{2}\rangle_{t}=\int_{0}^{\infty}r^{2}P_{t}(r)\,dr=4Dt.$ (3)
3.1.2 Empirical recapture ratios
@@
With the above context in place, we now introduce the four methodological approaches mentioned above, which can be applied to estimate the diffusion coefficient. Estimates obtained from the first three methods will be used as baselines for comparison with RDA-PSO estimates.
3.2 MDT-based model
We start with a practical estimate of the diffusion coefficient based on the mean distance traveled (MDT). As noted in section 2.3, the empirical MDT is often reported in the MRR literature, or can easily be estimated from those reports (see for instance [50]). Because results based on a few dozens of mosquitoes are unlikely to give reliable or consistent estimates from day to day, it is essential to aggregate mosquito counts obtained over different days in order to improve the estimate, since data from a larger number of walkers are then used to arrive at a single number. To this end, we define the MDT of an entire experiment, MDTtot, as a weighted average of the daily MDT’s, denoted by MDTi for day $i$, and apply the first equation of (3):
$MDT_{tot}$ $=\sum_{i=1}^{n}\tau_{i}\,MDT_{i}=\sum_{i=1}^{n}\tau_{i}\,\sqrt{\pi Dt_{i}}=\sqrt{\pi D}\sum_{i=1}^{n}\tau_{i}\sqrt{t_{i}},$ (6)
where $n$ is the number of days mosquitoes were collected, $\tau_{i}$ is the temporal ratio corresponding to day $i$ defined in (4) and $t_{i}$ is the number of days elapsed since release. We emphasize that in Equation (6), the quantity $MDT_{tot}$ defines an overall MDT, calculated for all recaptured mosquitoes, regardless of the day of recapture. This quantity is expected to be close to the empirical MDT, $MDT_{emp}$, reported in the literature (e.g. as defined in [41]). Indeed,
$\simeq\sum_{i=1}^{n}\tau_{i}\,MDT_{i}=MDT_{tot}.$
In the above, $N_{tot}=\sum_{i=1}^{n}N_{i}$ is the total number of recaptured mosquitoes, $N_{i}$ is the number of mosquitoes captured in day $i$, and $r_{k}$ is the distance from the release point where the $k$-th mosquito recaptured on day $i$ was caught. The difference between $MDT_{emp}$ and $MDT_{tot}$ is that the exact expression of $MDT_{i}$ is used in the latter, instead of its empirical approximation $\sum_{k=1}^{N_{i}}\frac{r_{k}}{N_{i}}$ used in the former. However, assuming that the two values are close, knowledge of the empirical MDT, along with equation (6), leads to the following estimate of the diffusion coefficient,
$D_{MDT}=\frac{1}{\pi}\left(\frac{MDT_{emp}}{\sum_{i=1}^{n}\tau_{i}\sqrt{t}_{i}}\right)^{2},$ (7)
where $\tau_{i}$ is the temporal ratio for day $i$ and $t_{i}$ is the number of days since release.
@@ 3.3 Time-corrected model
One of the main difficulties of translating MRR data into estimates of dispersal parameters is that only a very small percentage of released mosquitoes is eventually recaptured. To improve empirical estimates of the quantities defined in Section 3.1.1, it is therefore useful to combine information associated with different collection dates. To this end, we estimate the probability density function for a mosquito being present at a distance $r$ from the point of release at any time during the entire recapture experiment,
$P(r)=\int_{0}^{\infty}P_{t}(r)dt,$
as a weighted average of the probabilities $P_{t}(r)$ defined in Equation (1), where the weights reflect the fraction of mosquitoes recaptured each day $i$. Thus, we set $P(r)\simeq P_{tot}(r)$, with
@@
where $n$ is the total number of days mosquitoes were collected. To understand why weighting each term by $\tau_{i}$ is appropriate, first consider the case where there is no mosquito depletion. In such a situation, if we assume that recapture is proportional to presence, the number of mosquitoes captured on day $i$ does not depend on $t_{i}$, since $\int_{0}^{\infty}P_{t}(r)dr=1$ for all $t$. This implies that $\tau_{i}=1/n$. Consequently, the total number of mosquitoes collected over $n$ days is
$KN\int_{0}^{\infty}\left(\sum_{i=1}^{n}\tau_{i}\,P_{t_{i}}(r)\right)dr=KN\sum_{i=1}^{n}\left(\frac{1}{n}\int_{0}^{\infty}P_{t}(r)dr\right)=KN,$
where $N$ is the number of mosquitoes initially released and $K$ measures the total recapture ratio. However, in the presence of depletion, due to mosquitoes being caught, flying out of the study zone, or dying, the number captured on day $i$ is $KN_{i}$, where $N_{i}$ is close to but slightly less than $N$, which leads to $\tau_{i}=N_{i}/\left(\sum_{i=1}^{n}N_{i}\right)$. The definition of $P_{tot}$ is such that
$\int_{0}^{\infty}P_{tot}(r)dr=\int_{0}^{\infty}\left(\sum_{i=1}^{n}\frac{N_{i}}{\sum_{i=1}^{n}N_{i}}\,P_{t_{i}}(r)\right)dr=1,$
@@
Estimates of $Q_{tot}$ (and $P_{tot}$) may be compared directly to data observed in MRR experiments, and an estimate of the diffusion coefficient $D$ may be obtained by optimizing the fit between model and experiment. We therefore define the time-corrected model,
@@ 3.4 Area-and-time-corrected model
where $N_{j}$ is the number of mosquitoes actually captured in zone $j$, and $nT_{j}\neq 0$ is the number of traps placed in zone $j$. The above equation can be recast as $N_{j}^{c}=N_{j}\cdot\rho_{tot}/\rho_{j}$, with the symbol $\rho$ denoting the density of traps, which shows how the correction compensates for uneven trap density. Such a correction changes the number of mosquitoes recaptured in each zone every day of the experiment and thus affects the spatial and temporal ratios defined in Equations (5) and (4). Fitting these corrected ratios into (11) leads to a corrected estimate of the diffusion coefficient, which we refer to as $D_{ATC}$.